# Regimes and Transitions of the Nonlinear Temporal Talbot Effect: Underlying Mechanism for A-Type Breathers, Soliton Crystals, and Soliton Gas

### v1

@@ 3.1 Modelling of Light Propagation in Fibers
with $J_{k}(m)$ being modulation-depth-dependent Bessel functions of the first kind. It is a frequency comb centered at central frequency $\omega_{0}$ with line spacing $2\pi\Omega.$ With that, it satisfies the condition of the linear temporal Talbot effect [19] (Fig. 1, right).
@@ 3.2 Soliton Radiation Beat Analysis
From the implementation point of view, performing standard SRBA involves five steps [61], [63], [62].
  1. i)
    Numerical Simulation: Propagate an initial optical field $A(z=0,t)$ through an NLS (Eq. 2) with realistic parameters $\beta_{2}$ and $\gamma$. As the resolution of SRBA images scales with the fiber length, sufficiently long propagation distances $z$, typically several hundreds of kilometers, are required to obtain images with well-resolved spatial frequency structure. Therefore, I use a fiber length of $L=500~\text{km}$ for SRBA studies within the context of this study.
  2. ii)
@@ 4.1 Dispersion Relation of the Nonlinear Schrödinger Equation
$Z_{nm}=|\zeta_{n}(\omega_{n})-\zeta_{m}(\omega_{m})|.$ (12)
These quantities can be observed as fans, pitchfork-shaped, or parabolic ridges in the SRBA plots. Each field mode $\zeta_{0}(\omega_{0}),$ $\zeta_{\pm 1}(\omega_{\pm 1})$, $\zeta_{\pm 2}(\omega_{\pm 2})$ etc., gives rise to a separate ridge in the SRBA plot. The ridges move non-uniformly with $P_{0}$, depending on their Bessel function amplitudes and interaction terms, which results in an unequal spacing between them.
@@
Self-Phase Modulation (SPM): $\zeta_{k}(\omega_{k})\propto\gamma P_{0}J_{k}^{2}$ describes the nonlinear phase shift that the spatial mode $\zeta_{k}(\omega_{k})$ experiences due to its intensity ([66]). In the SRBA plot, it causes the ridge associated with $\omega_{k}$ to move to higher spatial frequencies $Z$ with increasing input power $P_{0}.$ In the temporal domain, the recurrence of pulse trains happens at shorter propagation distances as seen in Fig. 3, top (cf. [8], [60]).
Cross-Phase Modulation (XPM): $\zeta_{k}(\omega_{k})\propto 2\gamma P_{0}\sum_{n\neq k}J_{n}^{2}$ describes the nonlinear phase shift of the spatial mode $\zeta_{k}(\omega_{k})$ due to the influence of other frequency-comb components at $\omega_{n\neq k}$ ([66], [58]). As the comb line intensities are not homogeneous (cf. Fig. 1, right), XPM leads to non-uniform shifts in the spatial frequencies of different SRBA ridges. This, in turn, causes their curvature and changes their spacing relative to each other as the input power $P_{0}$ increases. Apart from that, XPM contributes to the broadening of the initial comb by changing the nonlinear phase shift that determines phase matching for FWM (cf. [59], [58], cf. [69]).
@@
In a regular FWM process, a nonlinear interaction between three spectral components $\omega_{n},$ $\omega_{m}$ and $\omega_{\ell}$ generates a fourth component, $\omega_{k}.$ These process yields a limited number of new frequency components, all depending on the number and phase-matching of input frequencies ([66]). A cascaded four-wave mixing is a sequential process where the FWM products generated by an initial FWM interaction act as inputs for further FWM processes. It yields many additional frequency components, allowing for the generation of broadband frequency combs through a cascade of FWM processes (cf. [70], [71], [72]).
For a satisfied phase-matching condition, the input power $P_{0}$ decides the type of FWM. Thus, higher input power enables the transition from regular to cascaded four-wave mixing by amplifying new frequencies to levels sufficient for further nonlinear interactions, favoring the cascaded process [73], [74]. As shown in Sec. 4.5, the PM modulation depth $m$ is another important parameter that decides the type of FWM.
@@
To conclude, the dispersion relation of spatial modes ($\zeta_{k}(\omega_{k})$ allows me to provide a physical interpretation of SRBA ridge patterns. It explains how GVD, SPM, XPM, and FWM modify the spatial modes, influencing SRBA ridge positions, curvature, and merging in the SRBA plots. Additionally, it highlights how increasing input power drives transitions from weakly interacting comb lines to phase-locked states, and ultimately to spatio-temporally localized solitons.
@@
Regime I and the Harmonics of the Linear Temporal Talbot Effect for $P_{0}\rightarrow 0~\text{W}.$ Here, we see the emergence of ridges at higher harmonics and lower harmonics of the fundamental Talbot frequency $Z_{T}$ for $P_{0}\rightarrow 0~\text{W}.$ The Talbot frequency relates to the linear temporal Talbot effect and constitutes the inverse of the Talbot period, i.e.
@@
The seeding of the fundamental Talbot frequency as an SRBA ridge is explicable by the beating of the spatial mode $\zeta_{0}$ of the central frequency $\omega_{0}$ with the neighboring modes $\zeta_{\pm 1}=\frac{\beta_{2}}{2}(2\pi\Omega)^{2}$ for $P_{0}\rightarrow 0~\text{W}$ (Eq. 9).
The appearance of higher and lower harmonics of $Z_{T}$ was already reported, but not exhaustively explained in Ref. [60]. To understand these harmonics of $Z_{T},$ let us consider two examples of Eq. 11 using only a few comb lines. Thus, a pair of lines $\omega_{k}$ and $\omega_{l}$ would produce the following beating pattern:
$P(z)=P_{0}|i^{k}J_{k}e^{i\zeta_{k}z}+i^{l}J_{l}e^{i\zeta_{l}z}|^{2}=P_{0}\left(J^{2}_{k}+J_{l}^{2}+2J_{k}J_{l}\cos{\left([\zeta_{k}-\zeta_{l}]z+\frac{\pi}{2}[k-l]\right)}\right)=$ (14)
$=P_{0}\left(J^{2}_{k}+J_{l}^{2}+2J_{k}J_{l}\cos{\left(\frac{Z_{T}}{2}[k^{2}-l^{2}]z+\frac{\pi}{2}[k-l]\right)}\right).$
Here, a pairwise beating between the lines with high values of the indices $k$ and $l$ can generate higher harmonics at multiples of $Z_{T}/2.$
@@
with an overlap of cosine functions of different periods being proportional to $\frac{Z_{T}}{2}[k^{2}-l^{2}],$ $\frac{Z_{T}}{2}[l^{2}-m^{2}],$ and $\frac{Z_{T}}{2}[k^{2}-m^{2}]$. Depending on the line-indices values, constructive overlap, i.e., interference between these cosines, can happen at longer propagation distances, resulting in subharmonics of the SRBA plot. Phase factors $\phi_{ij}:=\frac{\pi}{2}[i-j]$ play a role in achieving constructive interference of spatial modes or preventing it.
Already, these two simple examples of pairwise and 3-line (higher-order) interference show that the optical power $P(z)$ oscillates with fractional multiples of $Z_{T}/2$. This interference translates to the appearance of higher and lower harmonics of $Z_{T}$ in the SRBA plot for $P_{0}\rightarrow 0~\text{W}.$ The effect behind is the linear temporal Talbot effect [19], [20], [23]. As the richest and most interesting ridges arise at subharmonics of $Z_{T},$ I conclude that it is the higher-order interference (or beating) between several comb lines that sets the tone for the dynamics of a phase-modulated CW laser field (Fig. 2, left).
@@
Whereas the authors of Ref. [54] impose a phase shift of $\frac{\pi}{2}$ between the pump and the signal waves in their 3-wave experiment to generate doubly-periodic A-type solutions of the NLS, such a phase shift is naturally introduced by the linear temporal Talbot effect in a phase-modulated CW field [19], [20].
As seen in Fig. 2, left, A-type breather solutions are recognizable by pitchfork or rake-shaped SRBA ridge bundles. The most intense A-type breathers are seeded by subharmonics of the fundamental Talbot frequency $Z_{T},$ indicating higher-order interference between the comb lines and the corresponding spatial modes (Eqs. 9, 15) as a decisive mechanism for the formation of this type of NLS solutions.
@@
In Fig. 7, $P_{0}=0.27~\text{W}$ we see that solitons that evolved in regime III do not match the description of Peregrine solitons, as the latter sit on a homogeneous background ([80], [43], [44], [45], [81]). Here, however, the spatio-temporal localization goes almost down to zero, constituting soliton crystals as a spatio-temporally periodic soliton compound and a separate class of NLS solutions ([63], [64]). The regular spatio-temporal order is seeded by the higher-order interference of the temporal Talbot effect (cf. Eq. 15).
@@
The richness of harmonics above $Z_{T}$ is attributable to the beating of lines with high indices (cf. Eqs. 14, 15). Such frequencies are the result of cascaded FWM. In other words, soliton crystals in single-pass systems ([63], [64], [60]) result from cascaded FWM, contrary to A-type breathers that originate from regular FWM.
@@
At $P_{0}=0.27~\text{W}$, the spatial modes are (still) well phase-locked due to FWM of regime III, but significantly dephase due to SPM with increasing input power (Fig. 6, compare $P_{0}=0.27~\text{W}$ and $P_{0}=0.5~\text{W}).$ As a result, the rigidness of the spatio-temporal distribution initially imprinted by the Talbot effect dissolves, giving stage to the development of soliton gas, a state of several weakly interacting solitons (  [65]). Their irregular presence is well seen in the embedding of several homoclinic orbits within each other (Fig. 3, $P_{0}=0.5~\text{W}$).
@@
In conclusion, the analysis of input-power-dependent SRBA ridge evolution reveals that regime transitions in the temporal Talbot effect are governed by a shifting balance between GVD, SPM, XPM, and FWM in the spatial mode dispersion relation. At low powers, higher-order interference from the linear Talbot effect dominates, while increasing input power drives spectral broadening, spatial-mode phase locking, and spatio-temporal pulse localization through regular and cascaded FWM. Regular FWM leads to the formation of A-type breathers, whereas cascaded FWM involves the generation of soliton crystals. At the highest powers, SPM prevails, breaking spatial-mode phase locking and leading to a soliton gas of weakly interacting pulses.
@@ 4.4 Soliton Encoding by Frequency Comb Lines
Fig. 8 shows an example for the SRBA of the central comb line and lines with indices $k=2$ and $k=4.$ In this figure, green arrows indicate the most intense line-specific solitonic ridges. Additionally, we can observe black stripes whose input-power-dependent region increases with the index of the line. In the framework of SRBA, black stripes indicate that there is no oscillation over the propagation distance for the given line and input power; the amplitude of the line is too weak to participate in the nonlinear interaction with other lines via FWM and XPM. We also see that the regime of A-type breathers is covered by lines that are close to the central line (Line = 2, as an example). This observation supports the theory of regular FWM involving only a few lines in this regime. The regime of soliton crystals needs a higher number of comb lines to be set off (Line = 4). Therefore, the black stripe is broader. This observation supports the theory of cascaded FWM being primarily involved in the formation of these waves.
@@ 4.5 Effect of the Phase-Modulator Modulation Depth
Let us now study the impact of the PM modulation depth $m$ (Eq. 3) on the development of the nonlinear temporal Talbot effect regimes. Fig. 9, top, shows SRBA plots for different values of modulation depth, ranging from $m=0.1$ (weak modulation) to $m=1$ (quite strong modulation with dynamics studied in previous sections). The corresponding bottom panel shows how the number of initial comb lines of a phase-modulated CW-laser fields increases with the value of $m.$
@@
As shown in the next step, this observation supports the discussions in Sections 4.1 and 4.2, particularly the discussion about the impact of the type of FWM on the dynamics. That is, regular FWM with a low number of comb lines involved (initial and generated ones) leads to A-type breathers, whereas cascaded FWM yields soliton crystals. Considering the optical-power evolution and the trajectories for different modulation depths $m$ will help us understand different regimes in Fig. 9, top, and their smooth evolution from one to another with increasing $m.$
@@
To demonstrate that the dynamics is governed by separated (Peregrine) solitons rather than by collective states of breathers and crystals for low values of $m,$ let us consider the limit case of a weak modulation $m<<1.$ A Bessel function of the first kind that describes the comb-line amplitude (Eqs. 3, 5) can be written as
@@
for positive integer orders $n\geq 0.$ Taylor expansion of this function for $m<<1$ gives us following relationships for the central line at $\omega_{0}$ and the lines at $k=1$ and $k=2$ (cf. [86]):
@@
Here, we see that spatial modes of the most central lines collapse to a relationship between the GVD and self-phase modulation for $m\rightarrow 0$ (cf. Eq. 17), which is a characteristic of solitons ([66], [75]). It confirms the observation that weak modulation yields (Peregrine) soliton formation; collective states such as breathers and soliton crystals are not even possible. However, with increasing input power and the value of $m,$ the contributions of other nonlinear effects such as FWM and XPM become possible. In particular, FWM allows the creation of collective states of A-type breathers and soliton crystals (Figs. 10, 11).
@@
From the perspective of the temporal Talbot effect, an increasing number of initial comb-lines increases the order of interference of spatial field modes (Eq. 9), introducing SRBA ridges at multiples of $Z_{T}/2$ (Sec. 4.3). These (arising) number ridges correspond to spatial oscillations over shorter FPUT recurrence periods. From this observation and observations made in Sec. 4.3, I conclude that FPUT recurrence of breathers directly relates to the interference order induced by the temporal Talbot effect. More studies are certainly needed to better link FPUT recurrence with the temporal Talbot effect, thereby developing an exhaustive theory of the nonlinear temporal Talbot effect.

### v2

@@ 3.1 Modelling of Light Propagation in Fibers
with $J_{k}(m)$ being modulation-depth-dependent Bessel functions of the first kind. It is a frequency comb centered at central frequency $\omega_{0}$ with a line spacing of $2\pi\Omega.$ With that, it satisfies the condition of the linear temporal Talbot effect [19] (Fig. 1, right).
@@ 3.2 Soliton Radiation Beat Analysis
Performing standard SRBA involves five steps with theoretical background and a detailed implementation guide described in Refs. [61],[62], [63]. Here, I provide a summary needed to better undertand further stages of this study.
  1. 1.
    Numerical Simulation: Propagate an initial optical field $A(z=0,t)$ through an NLS (Eq. 2) with realistic parameters $\beta_{2}$ and $\gamma$. As the resolution of SRBA images scales with the fiber length, sufficiently long propagation distances $z$, typically several hundreds of kilometers, are required to obtain images with well-resolved spatial frequency structure. Throughout this study, I use a fiber length of $L=500~\text{km}$ for SRBA images. This value is a trade-off between the fiber length required for high-resolution SRBA images and computational time.
  2. 2.
@@ 4.1 Dispersion Relation of the Nonlinear Schrödinger Equation
$Z_{nm}=|\zeta_{n}(\omega_{n})-\zeta_{m}(\omega_{m})|$ (12)
with spatial modes $\zeta_{k}(\omega_{k})$ as given by the dispersion relation Eq. 9.
The quantities $Z_{nm}$ can be observed as fans, pitchfork-shaped, or parabolic ridges in the SRBA plots. Each field mode $\zeta_{0}(\omega_{0}),$ $\zeta_{\pm 1}(\omega_{\pm 1})$, $\zeta_{\pm 2}(\omega_{\pm 2})$ etc., gives rise to a separate ridge in the SRBA plot. The ridges move non-uniformly with $P_{0}$, depending on their Bessel function amplitudes and interaction terms, which results in an unequal spacing between them.
@@
Self-Phase Modulation (SPM): $\zeta_{k}(\omega_{k})\propto\gamma P_{0}J_{k}^{2}$ describes the nonlinear phase shift that the spatial mode $\zeta_{k}(\omega_{k})$ experiences due to its intensity ([66]). In the SRBA plot, it causes the ridge associated with $\omega_{k}$ to move to higher spatial frequencies $Z$ with increasing input power $P_{0}.$ In the temporal domain, it implies that the recurrence of pulse trains happens at shorter propagation distances as seen in Fig. 3, top (cf. [8], [60]).
Cross-Phase Modulation (XPM): $\zeta_{k}(\omega_{k})\propto 2\gamma P_{0}\sum_{n\neq k}J_{n}^{2}$ describes the nonlinear phase shift of the spatial mode $\zeta_{k}(\omega_{k})$ due to the influence of other frequency-comb components at $\omega_{n\neq k}$ ([66], [58]). As the comb line intensities are not homogeneous (cf. Fig. 1, right), XPM leads to non-uniform shifts in the spatial frequencies of different SRBA ridges. This, in turn, introduces a curvature of SRBA ridges and changes their spacing relative to each other as the input power $P_{0}$ increases. Apart from that, XPM contributes to the broadening of the initial comb by changing the nonlinear phase shift that determines phase matching for FWM (cf. [59], [58], cf. [69]).
@@
In a regular FWM process, a nonlinear interaction between three spectral components $\omega_{n},$ $\omega_{m}$ and $\omega_{\ell}$ generates a fourth component, $\omega_{k}.$ This process yields a limited number of new frequency components, all depending on the number and phase matching of input frequencies ([66]). A cascaded four-wave mixing is a sequential process where the FWM products generated by an initial FWM interaction act as inputs for further FWM processes. It yields many additional frequency components, allowing for the generation of broadband frequency combs through a cascade of FWM processes (cf. [70], [71], [72]). Please note that there is no strict definition of regular and cascaded FWM, and no clear difference between them is available in the literature. In its core, it is the same wave-mixing process governed by the Kerr nonlinearity for both types of FWM. The difference is rather made based on the role of generated waves: in cascaded FWM, newly generated comb lines serve as a source of further FWM products, whereas in regular FWM, they do not.
For a satisfied phase-matching condition, the input power $P_{0}$ decides over the type of FWM. Thus, higher input power enables the transition from regular to cascaded four-wave mixing by amplifying new frequencies to levels sufficient for further nonlinear interactions, favoring the cascaded process [73], [74]. As shown in Sec. 4.5, the PM modulation depth $m$ is another important parameter that is decisive for the type of FWM. The difference between the type of FWM becomes important in the differentiation between the A-type breathers and soliton crystals in Sec. 4.3.
@@
To conclude, the dispersion relation of spatial modes ($\zeta_{k}(\omega_{k})$ allows us to provide a physical interpretation of SRBA ridge patterns. It explains how GVD, SPM, XPM, and FWM modify the spatial modes, influencing SRBA ridge positions, curvature, and merging in the SRBA plots. Additionally, it highlights how increasing input power drives transitions from weakly interacting comb lines to phase-locked states, and ultimately to spatio-temporally localized solitons.
@@
Regime I and the Harmonics of the Linear Temporal Talbot Effect for $P_{0}\rightarrow 0~\text{W}.$ Here, we see the emergence of ridges at higher harmonics and lower harmonics of the fundamental Talbot frequency $Z_{T}$ for $P_{0}\rightarrow 0~\text{W}.$ The Talbot frequency relates to the linear temporal Talbot effect and constitutes the inverse of the Talbot period $L_{T}=\frac{2\pi}{(2\pi\Omega)^{2}|\beta_{2}|}$ [60], i.e.
@@
The seeding of the fundamental Talbot frequency as an SRBA ridge is explicable by the beating of the spatial mode $\zeta_{0}$ of the central frequency $\omega_{0}=0~\text{THz}$ with the neighboring modes $\zeta_{\pm 1}=\frac{\beta_{2}}{2}(2\pi\Omega)^{2}$ for $P_{0}\rightarrow 0~\text{W}$ (Eq. 9).
The appearance of higher and lower harmonics of $Z_{T}$ was already reported, but not exhaustively explained in Ref. [60]. To understand these harmonics, let us consider two examples of Eq. 11 using only a few comb lines. Thus, a pair of lines $\omega_{k}$ and $\omega_{l}$ would produce the following beating pattern:
$P(z)=$ $P_{0}|i^{k}J_{k}e^{i\zeta_{k}z}+i^{l}J_{l}e^{i\zeta_{l}z}|^{2}=$ (14)
$=$ $P_{0}\left(J^{2}_{k}+J_{l}^{2}+2J_{k}J_{l}\cos{\left([\zeta_{k}-\zeta_{l}]z+\frac{\pi}{2}[k-l]\right)}\right).$
For $P_{0}\rightarrow 0~\text{W},$ Eq. 14 simplifies to:
$P(z)=P_{0}\left(J^{2}_{k}+J_{l}^{2}+2J_{k}J_{l}\cos{\left(\frac{Z_{T}}{2}[k^{2}-l^{2}]z+\frac{\pi}{2}[k-l]\right)}\right),$ (15)
which shows that a pairwise beating between the lines with high values of the indices $k$ and $l$ can generate higher harmonics at multiples of $Z_{T}/2.$
@@
with an overlap of cosine functions of different periods. For $P_{0}\rightarrow 0~\text{W}$, these periods are proportional to $\frac{Z_{T}}{2}[k^{2}-l^{2}],$ $\frac{Z_{T}}{2}[l^{2}-m^{2}],$ and $\frac{Z_{T}}{2}[k^{2}-m^{2}]$. Depending on the line-indices values, constructive overlap, i.e., interference between these cosines, can happen at longer propagation distances, resulting in subharmonics of the SRBA plot. Phase factors $\phi_{ij}:=\frac{\pi}{2}[i-j]$ play a role in achieving constructive interference of spatial modes or preventing it.
These two simple examples of pairwise and 3-line (higher-order) interference show that the optical power $P(z)$ oscillates with fractional multiples of $Z_{T}/2$ for $P_{0}\rightarrow 0~\text{W}$. The effect behind is the linear temporal Talbot effect [19], [20], [23]. This interference translates to the appearance of higher and lower harmonics of $Z_{T}$ in the SRBA plot for $P_{0}\rightarrow 0~\text{W}.$ As the richest and most interesting ridges arise at subharmonics of $Z_{T},$ I conclude that it is the higher-order interference (or beating) between several comb lines that sets the tone for the dynamics of a phase-modulated CW laser field (Fig. 2, left). In other words, the linear temporal Talbot effect seeds the spatio-temporal distribution of optical pulses through interference of field’s spatial modes. With increasing input power and non-negligible nonlinear effects, the spatio-temporal distribution will be modified according to Eq. 9, connecting temporal Talbot effect with FPUT phenomena in optical fibers.
@@
Whereas the authors of Ref. [54] impose a phase shift of $\frac{\pi}{2}$ between the pump and the signal waves in their 3-wave experiment to generate doubly-periodic A-type breather solutions of the NLS, such a phase shift is naturally introduced by the linear temporal Talbot effect in a phase-modulated CW field propagating through an SMF [19], [20].
As seen in Fig. 2, left, A-type breather solutions are recognizable by pitchfork or rake-shaped SRBA ridge bundles. The most intense A-type breathers are seeded by subharmonics of the fundamental Talbot frequency $Z_{T},$ indicating higher-order interference between the comb lines and the corresponding spatial modes (Eqs. 9, 16) as a decisive mechanism for the formation of this type of NLS solutions.
@@
In Fig. 7, $P_{0}=0.27~\text{W}$ we see that solitons that evolved in regime III do not match the description of Peregrine solitons, as the latter sit on a homogeneous background ([80], [43], [44], [45], [81]). Here, however, the spatio-temporal localization goes almost down to zero, constituting soliton crystals as a spatio-temporally periodic soliton compounds and rather a separate class of NLS solutions ([63], [64]). The regular spatio-temporal order is seeded by the higher-order interference of the temporal Talbot effect (cf. Eq. 16).
@@
The richness of harmonics above $Z_{T}$ is attributable to the beating of lines with high indices (cf. Eqs. 14, 16). Such frequencies are the result of cascaded FWM. In other words, soliton crystals in single-pass systems ([63], [64], [60]) result from cascaded FWM, contrary to A-type breathers that originate from regular FWM.
@@
At $P_{0}=0.27~\text{W}$, the spatial modes are (still) well phase-locked due to FWM of regime III, but significantly dephase due to SPM with increasing input power (Fig. 6, compare $P_{0}=0.27~\text{W}$ and $P_{0}=0.5~\text{W}).$ As a result, the rigidness of the spatio-temporal distribution initially imprinted by the Talbot effect dissolves, giving stage to the development of soliton gas, a state of multiple weakly interacting solitons (  [65]). Their irregular spatio-temporal distribution is reflected in the embedding of several homoclinic orbits within each other (Fig. 3, $P_{0}=0.5~\text{W}$).
@@
To conclude, i) the initial spatio-temporal distribution of the optical pulses is seeded by (higher-order) interference of the spatial modes, which is due to the linear temporal Talbot effect; ii) the Kerr nonlinearity of the fiber and its effects modify this spatio-temporal distribution and can be integrated in the framework of the nonlinear temporal Talbot effect by considering frequency-dependent spatial modes. Regime transitions of the nonlinear temporal Talbot effect are governed by a shifting balance between GVD, SPM, XPM, and FWM in the spatial-mode dispersion relation. At low powers, higher-order interference from the linear Talbot effect dominates, while increasing input power drives spectral broadening, spatial-mode phase locking, and spatio-temporal pulse localization through regular and cascaded FWM. Regular FWM leads to the formation of A-type breathers, whereas cascaded FWM involves the generation of soliton crystals. At the highest powers, SPM prevails, breaking spatial-mode phase locking and leading to a soliton gas of weakly interacting pulses.
@@ 4.4 Soliton Encoding by Frequency Comb Lines
Fig. 8 shows an example for the SRBA of the central comb line and lines with indices $k=2$ and $k=4.$ In this figure, green arrows indicate the most intense line-specific solitonic ridges. Additionally, we observe black stripes whose input-power-dependent region increases with the index of the line. In the framework of SRBA, black stripes indicate that there is no oscillation over the propagation distance for the given line and input power; the amplitude of the line is too weak to participate in the nonlinear interaction with other lines via FWM and XPM. We also see that the regime of A-type breathers is covered by lines that are close to the central line (Line = 2). This observation supports the hypothesis of regular FWM involving only a few lines in this regime (Sec. 4.3). The regime of soliton crystals needs a higher number of comb lines to be set off (Line = 4). Therefore, the black stripe is broader. This observation supports the hypothesis of cascaded FWM being primarily involved in the formation of these waves.
@@ 4.5 Effect of the Phase-Modulator Modulation Depth
Let us now study the impact of the PM modulation depth $m$ on the development of the nonlinear temporal Talbot effect regimes. According to the Jacobi-Anger expansion of the initial condition, the parameter $m$ directly controls the number and relative amplitudes of frequency-comb lines through the Bessel-function coefficients $J_{k}(m)$ (Eqs. 3, 4). Thus, higher values of $m$ yield an initial comb with more lines. The aim of this section is to show that the PM modulation depth is an additional parameter in the dynamics governed by the nonlinear temporal Talbot effect. Other system paramters remain the same as in Sec. 2.
Fig. 9, top, shows SRBA plots for different values of modulation depth, ranging from $m=0.1$ (weak modulation) to $m=1$ (quite strong modulation with dynamics studied in previous sections). The corresponding bottom panel shows how the number of initial comb lines increases with the value of $m.$
@@
As shown in the next step, this observation supports the discussions in Sections 4.1 and 4.2, particularly the discussion about the impact of the type of FWM on the dynamics. That is, regular FWM with a low number of comb lines involved (initial and generated ones) leads to A-type breathers, whereas cascaded FWM yields soliton crystals. Considering the optical-power evolution and the trajectories for different modulation depths $m$ will help us understand different regimes in Fig. 9, top, and their evolution from one to another with increasing $m.$
@@
To demonstrate that the dynamics is governed by separated (Peregrine) solitons rather than by collective states of breathers and crystals for low values of $m,$ let us consider the limit case of a weak modulation $m<<1.$ A Bessel function of the first kind that describes the comb-line amplitude with index $n$ (Eqs. 3, 5) can be written as
@@
for positive integer orders $n\geq 0.$ Taylor expansion of this function for $m<<1$ gives us following relationships for the central line at $\omega_{0}$ and the lines at $k=1$ and $k=2$ next to it (cf. [86]):
@@
Here, we see that spatial modes of the most central lines collapse to a relationship between the GVD and self-phase modulation for $m\rightarrow 0$ (cf. Eq. 18), which is a characteristic of solitons ([66], [75]). The effects that build collective states, namely, FWM and to some degree XPM, are missing. It confirms the observation that weak modulation yields (Peregrine) soliton formation; collective states such as breathers and soliton crystals are not even possible. However, with increasing input power and the value of $m,$ the contributions of FWM and XPM become possible, leading to the formation of A-type breathers and soliton crystals (Figs. 10, 11).
@@
From the perspective of the temporal Talbot effect, an increasing number of initial comb-lines increases the order of interference of spatial field modes (Eq. 9), introducing SRBA ridges at multiples of $Z_{T}/2$ (Sec. 4.3). These (arising) ridges correspond to spatial oscillations over shorter FPUT recurrence periods. From this observation and observations made in Sec. 4.3, I conclude that FPUT recurrence of breathers directly relates to the interference order induced by the temporal Talbot effect. More studies are certainly needed to better link FPUT recurrence with the temporal Talbot effect, thereby developing an exhaustive theory of the nonlinear temporal Talbot effect.